A

child plays with a pair of scissors and

a piece of string 10 cm long. He cuts the string

into two at a randomly chosen place.

What is the probability that the piece of string to the left

of the pair of scissors is less than 4 cm long?



1
Expert's answer
2022-05-06T11:28:06-0400

Continuous uniform distribution

Let X be length of string to the left of the pair of scissors. "X \u223c U(0, 10)."

The density function of the continuous uniform random variable X on the

interval "[0, 10]" is


"f(x) = \begin{cases}
 \dfrac{1}{10} &0\le x\le10 \\
 0, &\text{elsewhere } 
\end{cases}"

"P(X<4)=\dfrac{4}{10}=\dfrac{2}{5}=0.4"